THE SUM OF TOM'S, DICK'S, HARRY'S AGES IS 31. TOM IS TWICE AS OLD AS DICK. HARRY IS 6 YEARS OLDER THAN TOM. USE ALEGBRA TO FIND EACH MAN'S AGE.
Answer by Alan3354(69443)

THE SUM OF TOM'S, DICK'S, HARRY'S AGES IS 31. TOM IS TWICE AS OLD AS DICK. HARRY IS 6 YEARS OLDER THAN TOM. USE ALEGBRA TO FIND EACH MAN'S AGE.
-------------
t + d + h = 31
t = 2d
h = t + 6
---------
Sub for t
2d + d + h = 31
3d + h = 31
h = 2d + 6
-----------
Sub for h
3d + 2d+6 = 31
5d = 25
d = 5
------
t = 10
h = 16